### v3

@@ 6 The Pressure Vessel Design Problem
Figure 14: The Structure of the Vessel [14]
Figure 15: The Heatmaps of the Vessel [18]
@@
As Table 7 and Fig.16 illustrated, the HEO algorithm show a relatively well small cost among the others, but larger than the cost of QPSO. The Pressure Vessel Design Problem is a linear combination problem with constraints, which is similar to the case of the Step Function. From the HEO’s cost in the Step Function(or $F_{2}$) being larger than QPSO,it is not unexpected that HEO perform worse than QPSO in this problem, which might indicate HEO’s weakness when solving the complex problem with simple constraints.
Figure 16: The Distribution of Cost for Pressute Vessel Design
7 Tubular Column Design Problem
As Fig.17 shown, the Tubular Column this problem is structural part that is a cylindrical shaft, which may be solid, or made of a single circular cross-section and produced from metal, concrete and other materials [1]. It is commonly used within constructing projects where it is required to support beams, as well as other elements, such as columns, and so on, and it is also used in bridges and other types of structures.
@@ 7 Tubular Column Design Problem
In Table 8, the paper set 5 algorithms with same parameter as the previous tests in the experiment(HEO with $c_{m}ax=1$), the AFSA is do not show any inclination of converge for testing engineering problems, hence it get rid from it. The HEO has the smallest mean value of cost compared to other algorithms in 30 times validations, might illustrate its good performance in the tubular column problem,or relatively simple function with complex constraints. As for the deviations, the GA get the smallest one,indicating its stability to some extent.
Figure 17: The Structure of the Tubular Column Design[1]
@@
8 Discussion
The experimental results indicate that HEO outperforms other algorithms in terms of convergence speed and solution quality across a range of benchmark functions. The algorithm’s ability to balance exploration and exploitation, coupled with its adaptive strategies, has resulted in competitive performance in challenging optimization landscapes. The TSP and test in Tubular Column Design also validates its availability in path-planning applications.
The discussion of TSP highlights the effectiveness of the HEO algorithm in general Path Planning Problems, as verified through the comparison with Tabu Search and Random Search methodologies. When dealing with a relatively complex situation, the findings indicate that HEO offers acceptable performance and potential, particularly in terms of searching speed.
@@ 9 Conclusion
Moreover, the adaptability of the HEO algorithm allows it to be applied to dynamic environments where the objective functions or constraints may change over time. By continuously adapting its search strategy and updating the particle positions, HEO can effectively track and respond to changing optimization landscapes, ensuring that high-quality solutions are maintained even in dynamic scenarios, including various engineering applications.
Overall, the HEO has demonstrated its effectiveness and adaptability in addressing complex optimization challenges. Future work could focus on further refining the algorithm’s parameters and exploring its application in practical optimization problems. Additionally, the algorithm could be extended to address multi-objective optimization problems and dynamic environments or change the strategy of escape and skip, further enhancing its versatility and applicability.
@@
The experiment is data-available for replications, provide the source code of HEO in: https://github.com/Spike8086/HEO.git, provide data of result and the notebook of experiment in: https://www.kaggle.com/code/spike8086/heo-formal-ver, and give the code for TSP test:https://www.kaggle.com/code/spike8086/heo-for-tsp

### v4

@@ 5.1 The Pressure Vessel Design Problem
Figure 13: The Structure of the Vessel-[11]
Figure 14: The Heatmaps of the Vessel-[14]
@@
As Table 6 and Fig.15 illustrated, the HEO algorithm show a relativelyx small cost among the others, but larger than the cost of QPSO. The Pressure Vessel Design Problem is a linear combination problem with constraints, which is similar to the case of the Step Function. From the HEO’s cost in the Step Function(or $F_{2}$) being larger than QPSO,it is not unexpected that HEO perform worse than QPSO in this problem, which might indicate HEO’s weakness when solving the complex problem with simple constraints.
Figure 15: The Distribution of Cost for Pressure Vessel Design
5.2 Tubular Column Design Problem
As Fig.16 shown, the Tubular Column this problem is structural part that is a cylindrical shaft, which may be solid, or made of a single circular cross-section and produced from metal, concrete, and other materials [1]. It is commonly used within constructing projects where it is required to support beams, as well as other elements, such as columns, and so on, and it is also used in bridges and other types of structures.
@@ 5.2 Tubular Column Design Problem
In Table 7, the paper sets 5 algorithms with the same parameter as the previous tests in the experiment(HEO with $c_{m}ax=1$), the AFSA does not show any inclination to converge to test engineering problems; hence, relevant data are eliminated in the table. The HEO has the smallest mean value of cost compared to other algorithms in 30 times validations, might illustrate its good performance in the tubular column problem,or relatively simple function with complex constraints. As for the deviations, the GA get the smallest one,indicating its stability to some extent.
Figure 16: The Structure of the Tubular Column Design-[1]
@@
6 Model-based Optimization for Real-time Problems
The experimental results indicate that HEO outperforms other algorithms in terms of convergence speed and solution quality across a range of benchmark functions. The algorithm’s ability to balance exploration and exploitation, coupled with its adaptive strategies, has resulted in competitive performance in challenging optimization landscapes. The discussion of stimulated engineering problems like Tubular Column Design Problem highlights the effectiveness of the HEO algorithm in general applications.
[…]
For further testing of the HEO, the paper would use the tested optimization algorithms and Grid Search(because it is a good baseline for hyperparameter tuning[12]) with 50 population size and 50 iterations(2500 size sample space for Grid Search) to optimize the Logistic Regression parameters for better performance in this classification data set, due to the Logistic Regression do not have any randomness(The definition of the Logistic Regression had shown in equation (6)). The first parameter is the Cost Parameter C in Logistic Regression, and the second one is the maximum iteration $i_{max}$ to minimize the mean square error shown in equation (21) of the model. This method that find candidate solutions for model’s parameter directly called the Model-based Optimization[3], Model-based Optimization provides a controlled and repeatable framework for testing, ensuring consistent and reliable results.
$h_{\theta}(x)$ $=\frac{1}{1+\exp{(-\theta^{T}x)}}$
$J(\theta)$ $=-\frac{1}{m}\sum_{i=1}^{m}[y^{(i)}\ln{(h_{\theta}(x^{(i)}))}+(1-y^{(i)})\ln{(1-h_{\theta}(x^{(i)}))}]$
$\frac{\partial J(\theta)}{\partial\theta_{j}}$ $=\frac{1}{m}\sum_{i=1}^{m}(h_{\theta}(x^{(i)})-y^{(i)})x_{j}^{(i)}$ (20)
$Consider:$
$\vec{x}=[x_{1},x_{2}]$
$Minimize:$
$f(\vec{x})=\sum\frac{((\beta_{1}x_{1}+\beta_{2}x_{2})-y)^{2}}{2}$
$Variable\quad Range:$
$1e-16\leq x_{1}\leq 100$
$1\leq x_{2}\leq 100$
6.1 Osmancik-97 and Cammeo Rice Classification
Rice industry is world-wide and have a demand to selecting out the ones that satisfy the various quality criteria of countries. The paper by Kay CINAR and Murat KOKLU suggest a new tabular dataset extracted from image characteristics to identify two rice species Osmancik-97 and Cammeo [4],a binary classification contains 2180 samples of Osmancik and 1630 for Cammeo. The performance of the models shown in the Table 8, with the same ratio(test set:train set is 0.25:0.75), the best model with HEO has a performance(accuracy: 0.9328, f1-score: 0.9328) that exceeds the record (accuracy: 0.9302, f1-score: 0.9180) in the original paper.The Evaluation Metrics including:
Accuracy $=\frac{TP+TN}{TP+TN+FP+FN}$ (22)
Sensitivity $=\frac{TP}{TP+FN}$
Specificity $=\frac{TN}{TN+FP}$
Precision $=\frac{TP}{TP+FP}$
Recall $=\frac{TP}{TP+FN}$
F1-score $=2\cdot\frac{Precision\cdot Recall}{Precision+Recall}$
Where: - $TP$ represents the number of true positives. - $TN$ represents the number of true negatives. - $FP$ represents the number of false positives. - $FN$ represents the number of false negatives.
Figure 17: The raw sample of the Rice Dataset-[4]
Table 8: Rice Classification in Test Set
PSO GWO HEO GA QPSO GridSearch
Accuracy 0.9317 0.9296 0.9328 0.9317 0.9296 0.9296
Sensitivity 0.9418 0.9381 0.9418 0.9418 0.9381 0.9381
Specificity 0.9181 0.9181 0.9205 0.9181 0.9181 0.9181
Precision 0.9317 0.9296 0.9328 0.9317 0.9296 0.9296
Recall 0.9317 0.9296 0.9328 0.9317 0.9296 0.9296
F1-score 0.9317 0.9296 0.9328 0.9317 0.9296 0.9296
SearchTime(s/it) 119.76 138.67 79.32 140.98 120.29 22.28
$x_{1}(C)$ 15.8281 11.6165 7.2574 7.6160 15.2641 6.0000
$x_{2}(i_{max})$ 60 74 77 62 74 51
@@ 7 Conclusion
Moreover, the adaptability of the HEO algorithm allows it to be applied to dynamic environments where the objective functions or constraints may change over time. By continuously adapting its search strategy and updating the position of particles, HEO can effectively track and respond to changing optimization landscapes, ensuring that high quality solutions are maintained even in dynamic scenarios, including various engineering applications.
Overall, the HEO has demonstrated its stable effectiveness and adaptability in addressing general optimization challenges. Future work could focus on further refining the algorithm’s parameters and exploring its application in practical optimization problems. Additionally, the algorithm could be extended to address multi-objective optimization problems and dynamic environments or change the strategy of escape and skip, further enhancing its versatility and applicability.
@@
The experiment is data-available for replications, provide the source code of HEO in: https://github.com/Spike8086/HEO.git, provide the data of the result and the notebook of the experiment in: https://www.kaggle.com/code/spike8086/heo-formal-ver